# Physics-Aware RIS Codebook Compilation for Near-Field Beam Focusing under Mutual Coupling and Specular Reflections

### v1

@@ Abstract
Next-generation wireless networks are envisioned to achieve reliable, low-latency connectivity within environments characterized by strong multipath and severe channel variability. Programmable wireless environments (PWEs) address this challenge by enabling deterministic control of electromagnetic (EM) propagation through software-defined reconfigurable intelligent surfaces (RISs). However, effectively configuring RISs in real time remains a major bottleneck, particularly under near-field conditions where mutual coupling and specular reflections alter the intended response. To overcome this limitation, this paper introduces MATCH, a physics-based codebook compilation algorithm that explicitly accounts for the EM coupling among RIS unit cells and the reflective interactions with surrounding structures, ensuring that the resulting codebooks remain consistent with the physical characteristics of the environment. Finally, MATCH is evaluated under a full-wave simulation framework incorporating mutual coupling and secondary reflections, demonstrating its ability to concentrate scattered energy within the focal region, confirming that physics-consistent, codebook-based optimization constitutes an effective approach for practical and efficient RIS configuration.
@@ I Introduction
While PWEs enable the dynamic control of EM propagation, the real-time configuration requirement imposes a fundamental need for low-complexity methodologies capable of coordinating thousands of unit cells simultaneously. To alleviate this computational burden, codebook-based solutions have been introduced, where a finite set of optimized reflection patterns is precomputed offline and stored for rapid deployment during operation [6, 5, 10, 7]. In more detail, codebook-based schemes eliminate the need for continuous channel estimation and iterative optimization, thereby ensuring timely adaptation to varying propagation conditions. However, most existing approaches have been developed under simplified far-field assumptions, which fail to capture the EM behavior that arises in the near-field regime, where mutual coupling among RIS unit cells and specular reflections from surrounding structures significantly influence the RIS response. In this direction, recent studies have investigated near-field codebook design [4] and coupling-aware modeling [11], yet these efforts generally treat the two effects in isolation or through approximate formulations.
In this paper, we present MATCH, a physics-based codebook compilation algorithm tailored for realistic RIS-assisted PWEs. MATCH targets near-field beam focusing and bridges geometric-optics initialization with physics-aware, multi-stage refinement. It exploits the physics knowledge gathered during its execution to accelerate convergence and manages the propagated field holistically to concentrate energy within the receiver’s region for enhanced performance. The performance of MATCH is assessed through full-wave simulations that capture complete EM interactions, including effects critical to the macroscopic RIS response such as mutual coupling among elements and specular reflections.
The remainder of the paper is organized as follows. Section II presents the EM modeling of the RIS-enabled environment. Section III details the workflow of the proposed MATCH algorithm. Section IV discusses the evaluation methodology and results, while Section V concludes the paper.
@@ III MATCH: A Novel Physics-Based Codebook Compilation Algorithm
To formulate the design objective of MATCH, the desired EM behavior is defined over specific regions of the environment where field enhancement or suppression is required. The receiver area corresponds to a focus region $\mathcal{F}$ centered at one or more points $\{\mathbf{c}^{(\ell)}\}{\ell=1}^{L}$, each modeled as a small sphere of radius $r$ that represents the spatial extent within which a strong and uniform field distribution is desired. To complement this focusing objective, a second region $\mathcal{O}$ is defined to represent the remaining part of the environment, excluding all focus spheres, where the field intensity should be minimized. However, since the EM field distribution is inherently coupled across space, the concentration of energy within $\mathcal{F}$ cannot be achieved independently of how the field behaves elsewhere. Consequently, an exclusive maximization of the field inside the target region may not effectively redirect the scattered energy toward it, as undesired constructive interference outside $\mathcal{F}$ could still dominate the overall response. A formulation that simultaneously enhances the field within $\mathcal{F}$ and attenuates it across $\mathcal{O}$ therefore ensures that the optimization acts coherently on the global energy balance of the environment, steering the configuration toward physically consistent focusing behavior.
To evaluate these objectives efficiently, the entire space is discretized into observation points at which $E(\mathbf{r})$ is computed, with a total of $N_{\mathcal{F}}$ samples distributed within the focus regions and $N_{\mathcal{O}}$ samples placed throughout the outer region. The sampling density is intentionally increased inside $\mathcal{F}$ to capture fine spatial variations of the field, whereas a coarser grid is used in $\mathcal{O}$ to maintain computational tractability. Based on this discretization, the first design goal of MATCH is to maximize the mean field magnitude $\bar{E}_{\mathcal{F}}$ within the focus regions, expressed as
@@
Following the GO initialization, MATCH transitions to a sensitivity-based local refinement stage (Stage 1) that builds directly upon the physics of the full propagation model in (5). At this stage, mutual coupling among unit cells and secondary wall reflections are incorporated to account for the complete EM interactions that characterize the RIS-assisted channel. To consistently integrate these effects within the optimization process, each RIS element undergoes small iterative phase perturbations of magnitude $\delta\phi$, and after each adjustment, the total field $E(\mathbf{r})$ is recomputed across the sampling sets to evaluate the corresponding variation in the focusing objective $\bar{E}_{\mathcal{F}}$. By continuously comparing successive field realizations, the procedure selectively retains phase updates that enhance the objective while reverting those that degrade it, thereby steering the configuration toward a physically consistent local optimum. Through this mechanism, the refinement process effectively performs a gradient-driven adaptation of the RIS phases, following the established principles of beamforming optimization, suitably extended to the near-field and strongly coupled regime of indoor RIS environments [1, 2]. The iterations proceed until the relative improvement in $\bar{E}{\mathcal{F}}$ falls below a local tolerance $\varepsilon_{\mathrm{loc}}$, indicating convergence to a locally stable configuration. In parallel, the same perturbation–response mechanism quantifies the contribution of each RIS element to the overall focusing performance through its sensitivity
@@ V Conclusion
In this paper, we introduced MATCH, a physics-based codebook compilation algorithm for RIS-assisted programmable wireless environments that achieves accurate near-field beam focusing by jointly accounting for mutual coupling and specular reflections. MATCH progresses from a geometric-optics initialization that establishes a coherent phase baseline, to a sensitivity-based refinement that incorporates full EM interactions, followed by a global exploration guided by element influence to target high-impact regions, and a final sensitivity-based refinement that consolidates convergence and stabilizes the near-field focus. Simulation results showed that coupling and reflections can be cooperatively exploited to direct nearly all scattered energy toward the focal region, achieving strong concentration and minimal leakage. Overall, MATCH exploits the physics knowledge gathered during its execution to accelerate convergence and manages the propagated field holistically, concentrating energy within the receiver’s region for enhanced performance. These results confirm that a physics-consistent, codebook-based formulation provides an effective and interpretable path for RIS optimization, linking practical focusing performance with clear physical insight into near-field energy redistribution.
@@ Acknowledgments
This work has received funding from the SNS-JU under the European Union’s Horizon Europe research and innovation programme, in the frame of the NATWORK project (No 101139285) and CYBERSECDOME (No 101120779).

### v2

@@ Abstract
Next-generation wireless networks are envisioned to achieve reliable, low-latency connectivity within environments characterized by strong multipath and severe channel variability. Programmable wireless environments (PWEs) address this challenge by enabling deterministic control of electromagnetic (EM) propagation through software-defined reconfigurable intelligent surfaces (RISs). However, effectively configuring RISs in real time remains a major bottleneck, particularly under near-field conditions where mutual coupling and specular reflections alter the intended response. To overcome this limitation, this paper introduces MATCH, a physics-based codebook compilation algorithm that explicitly accounts for the EM coupling among RIS unit cells and the reflective interactions with surrounding structures, ensuring that the resulting codebooks remain consistent with the physical characteristics of the environment. Finally, MATCH is evaluated under a realistic EM framework incorporating mutual coupling and secondary reflections, demonstrating its ability to concentrate scattered energy within the focal region, confirming that physics-consistent, codebook-based optimization constitutes an effective approach for practical and efficient RIS configuration.
@@ I Introduction
While PWEs enable the dynamic control of EM propagation, the real-time configuration requirement imposes a fundamental need for low-complexity methodologies capable of coordinating thousands of unit cells simultaneously. To alleviate this computational burden, codebook-based solutions have been introduced, where a finite set of optimized reflection patterns is precomputed and stored for rapid deployment during operation [5, 6]. In more detail, the codebook is compiled offline for a known and static deployment topology: for each supported functionality and candidate user/target location, an optimization loop computes the corresponding RIS reflection coefficients and stores the optimum as a codebook entry [4]. Since this procedure does not occur during operation, its computational load is not the limiting factor. In the research bibliography, most existing approaches have been developed under simplified far-field assumptions, which fail to capture the EM behavior that arises in the near-field regime, where mutual coupling among RIS unit cells and specular reflections from surrounding structures significantly influence the RIS response. In this direction, recent studies have investigated near-field codebook design [3] and coupling-aware modeling [9], yet these efforts generally treat the two effects in isolation or through approximate formulations.
In this paper, we present MATCH, a physics-based codebook compilation algorithm tailored for realistic RIS-assisted PWEs. MATCH targets near-field beam focusing and bridges geometric-optics initialization with physics-aware, multi-stage refinement. It exploits the physics knowledge gathered during its execution to accelerate convergence and manages the propagated field holistically to concentrate energy within the receiver’s region for enhanced performance. The performance of MATCH is assessed through physics-based EM simulations that capture complete EM interactions, including effects critical to the macroscopic RIS response such as mutual coupling among elements and specular reflections.
@@ II System Model
We model the RIS as ideal and lossless, using unit-modulus reflection coefficients. The framework directly extends to lossy RISs by imposing an amplitude constraint $|\Gamma_{n}|<1$ and/or using per–unit-cell measured (or simulated) amplitude–phase responses. We further assume identical RIS units with the same coupling characteristics, and account for mutual coupling via an effective parameter that mainly captures interactions within a local neighborhood of adjacent elements. This coupling model—its strength and neighborhood definition—can be re-parameterized for different RIS designs (e.g., element geometry, inter-element spacing, substrate, biasing network) using calibration measurements or manufacturer characterization.
@@ III MATCH: A Novel Physics-Based Codebook Compilation Algorithm
To formulate the design objective of MATCH, the desired EM behavior is defined over specific regions of the environment where field enhancement or suppression is required. The receiver area corresponds to a focus region $\mathcal{F}$ centered at one or more points $\{\mathbf{c}^{(\ell)}\}{\ell=1}^{L}$, each modeled as a small sphere of radius $r$ that represents the spatial extent within which a strong and uniform field distribution is desired. To complement this focusing objective, a second region $\mathcal{O}$ is defined to represent the remaining part of the environment, excluding all focus spheres, where the field intensity should be minimized. However, since the EM field distribution is inherently coupled across space, the concentration of energy within $\mathcal{F}$ cannot be achieved independently of how the field behaves elsewhere. Consequently, an exclusive maximization of the field inside the target region may not effectively redirect the scattered energy toward it, as undesired constructive interference outside $\mathcal{F}$ could still dominate the overall response. A formulation that simultaneously enhances the field within $\mathcal{F}$ and attenuates it across $\mathcal{O}$ therefore ensures that the optimization acts coherently on the global energy balance of the environment, steering the configuration toward physically consistent focusing behavior. To evaluate these objectives efficiently, the entire space is discretized into observation points at which $E(\mathbf{r})$ is computed, with a total of $N_{\mathcal{F}}$ samples distributed within the focus regions and $N_{\mathcal{O}}$ samples placed throughout the outer region. The sampling density is intentionally increased inside $\mathcal{F}$ to capture fine spatial variations of the field, whereas a coarser grid is used in $\mathcal{O}$ to maintain computational tractability. Based on this discretization, the first design goal of MATCH is to maximize the mean field magnitude $\bar{E}_{\mathcal{F}}$ within the focus regions, expressed as
@@
Following the GO initialization, MATCH transitions to a sensitivity-based local refinement stage (Stage 1) that builds directly upon the physics of the full propagation model in Eq. (5). At this stage, mutual coupling among unit cells and secondary wall reflections are incorporated to account for the complete EM interactions that characterize the RIS-assisted channel. To consistently integrate these effects within the optimization process, each RIS element undergoes small iterative phase perturbations of magnitude $\delta\phi$, and after each adjustment, the total field $E(\mathbf{r})$ is recomputed across the sampling sets to evaluate the corresponding variation in the focusing objective $\bar{E}_{\mathcal{F}}$. By continuously comparing successive field realizations, the procedure selectively retains phase updates that enhance the objective while reverting those that degrade it, thereby steering the configuration toward a physically consistent local optimum. Through this mechanism, the refinement process effectively performs a gradient-driven adaptation of the RIS phases, following the established principles of beamforming optimization, suitably extended to the near-field and strongly coupled regime of indoor RIS environments [1]. The iterations proceed until the relative improvement in $\bar{E}{\mathcal{F}}$ falls below a local tolerance $\varepsilon_{\mathrm{loc}}$, indicating convergence to a locally stable configuration. In parallel, the same perturbation–response mechanism quantifies the contribution of each RIS element to the overall focusing performance through its sensitivity
@@ V Conclusion
In this paper, we introduced MATCH, a physics-based codebook compilation algorithm for RIS-assisted programmable wireless environments that achieves accurate near-field beam focusing by jointly accounting for mutual coupling and specular reflections. MATCH progresses from a geometric-optics initialization that establishes a coherent phase baseline, to a sensitivity-based refinement that incorporates full EM interactions, followed by a global exploration guided by element influence to target high-impact regions, and a final sensitivity-based refinement that consolidates convergence and stabilizes the near-field focus. Overall, MATCH exploits the physics knowledge gathered during its execution to accelerate convergence and manages the propagated field holistically, concentrating energy within the receiver’s region for enhanced performance. These results confirm that a physics-consistent, codebook-based formulation provides an effective and interpretable path for RIS optimization, linking practical focusing performance with clear physical insight into near-field energy redistribution.
@@ Acknowledgments
This work has been funded by the Smart Networks and Services Joint Undertaking (SNS JU) under the European Union’s Horizon Europe research and innovation programme through the 6G Trans-Continental Edge Learning (6G-XCEL) project (Grant Agreement No 101139194), and through the NATWORK project (Net-Zero self-adaptive activation of distributed self-resilient augmented services; Grant Agreement No 101139285). Within 6G-XCEL, we established the initial technical foundations and performed the final synthesis and interpretation of results, while, within NATWORK, we developed the experimental workflow and carried out the main evaluation/validation.